Bind file: bind/emacs.bind

Key Description

C-˜ S-at  mark-on
C-a  line-begin
C-b  char-backward
C-d  delete-forward
C-e  line-end
C-f  char-forward
C-g  cancel
C-h  hyphenation-point-insert
C-i  hfill-insert
C-j  break-paragraph
C-k  line-delete-forward
C-l  screen-recenter
C-m  mark-toggle
C-n  down
C-o  inset-toggle

C-p up

C-s  find-replace
M-˜ S-percent find-replace
C-t  chars-transpose
C-u  font-underline
C-v  screen-down
C-w  cut
C-y  paste
C-S-Y  layout-paste

C-˜ S-slash undo C-˜ S-underscore undo

C-x a  buffer-auto-save
C-x d  buffer-new
C-x g  buffer-view ps
C-x k  buffer-close
C-x p  buffer-view dvi
C-x r  buffer-update dvi
C-x t  buffer-update dvi
C-x u  undo
C-x v h  vc-history
C-x v v  vc-check-in
C-x v u  vc-revert
C-x v c  vc-undo-last
C-x v i  vc-register

C-x bracketleft screen-up C-x bracketright screen-down

C-x C-a  buffer-auto-save
C-x C-b  menu-open documents
C-x C-c  lyx-quit
C-x C-d  buffer-new
C-x C-f  file-open
C-x C-g  buffer-view ps
C-x C-p  buffer-view dvi
C-x C-q  buffer-toggle-read-only
C-x C-s  buffer-write
C-x C-t  buffer-update dvi
C-x C-w  buffer-write-as
M-x  command-execute
C-Right  word-forward

                                                    C-Left  word-backward
                                                    C-Up  paragraph-up
                                                    C-Down  paragraph-down
                                                    Home  buffer-begin
                                                    End  buffer-end
                                                    C-˜ S-greater  reference-goto
                                                    C-˜ S-less  bookmark-goto 0
                                                    C-˜ S-1  bookmark-goto 1
                                                    C-˜ S-2  bookmark-goto 2
                                                    C-˜ S-3  bookmark-goto 3
                                                    C-M-˜ S-1  bookmark-save 1
                                                    C-M-˜ S-2  bookmark-save 2
                                                    C-M-˜ S-3  bookmark-save 3
                                                    S-Right  forward-select
                                                    S-Left  backward-select
                                                    S-C-F  forward-select
                                                    S-C-B  backward-select
                                                    S-Up  up-select
                                                    S-Down  down-select
                                                    S-C-Right  word-forward-select
                                                    S-C-Left  word-backward-select
                                                    S-C-A  line-begin-select
                                                    S-C-E  line-end-select
                                                    S-C-Up  paragraph-up-select
                                                    S-C-Down  paragraph-down-select
                                                    S-Home  buffer-begin-select
                                                    S-End  buffer-end-select
                                                    S-Prior  screen-up-select
                                                    S-Next  screen-down-select
                                                    S-C-Home  buffer-begin-select
                                                    S-C-End  buffer-end-select
                                                    S-KP_Right  forward-select
                                                    S-KP_Left  backward-select
                                                    S-KP_Up  up-select
                                                    S-KP_Down  down-select
                                                    S-C-KP_Right word-forward-select
                                                    S-C-KP_Left  word-backward-select
                                                    S-C-KP_Up  paragraph-up-select
                                                    S-C-KP_Down paragraph-down-select
                                                    S-KP_Home  line-begin-select
                                                    S-KP_End  line-end-select
                                                    S-KP_Prior  screen-up-select
                                                    S-KP_Next  screen-down-select
                                                    S-C-KP_Home buffer-begin-select
                                                    S-C-KP_End  buffer-end-select
                                                    S-C-KP_Prior buffer-begin-select
                                                    S-C-KP_Next buffer-end-select

C-KP_Insert copy S-KP_Insert paste S-KP_Delete cut

                                                    C-Delete  word-delete-forward
                                                    M-d  word-delete-forward
                                                    C-BackSpace  word-delete-backward
                                                    M-Return  break-paragraph-keep-layout

C-Return  break-line
C-S-L  ligature-break-insert
C-space  protected-space-insert
C-period  end-of-sentence-period-insert
M-period  dots-insert
M-w  copy
M-S-W  layout-copy
Escape  meta-prefix
Include file  menus.bind

Include file math.bind

Include file  latinkeys.bind
Include file  cyrkeys.bind
Include file  greekkeys.bind